A ladder is leaning against a building. The distance from the bottom of the ladder to the building is 8 ft less than the length of the ladder. How high up the building is the top of the ladder if that distance is 1 ft less than the length of the ladder.
Answer by stanbon(75887)

A ladder is leaning against a building. The distance from the bottom of the ladder to the building is 8 ft less than the length of the ladder. How high up the building is the top of the ladder if that distance is 1 ft less than the length of the ladder.
-----
Sketch the picture.
You have a right triangle
base = x-8
hypotenuse ladder = x
height = x-1
----------------
Equation::
x^2 = (x-8)^2 + (x-1)^2
------
x^2 = x^2 - 16x + 64 + x^2 - 2x + 1
------
x^2 = 2x^2 - 18x +65
----
x^2 -18x+65 = 0
------
x = 13 or x = 5
---
Acceptable answer: x = 13
Ans: height = x-1 = 12
------------
Cheers,
Stan H.